# MER 2024: Semi-Supervised Learning, Noise Robustness, and Open-Vocabulary Multimodal Emotion Recognition

### v1

@@ Abstract.
Multimodal emotion recognition is an important research topic in artificial intelligence. Over the past few decades, researchers have made remarkable progress by increasing dataset size and building more effective architectures. However, due to various reasons (such as complex environments and inaccurate labels), current systems still cannot meet the demands of practical applications. Therefore, we plan to organize a series of challenges around emotion recognition to further promote the development of this field. Last year, we launched MER2023111http://merchallenge.cn/mer2023, focusing on three topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue to organize MER2024222https://zeroqiaoba.github.io/MER2024-website. In addition to expanding the dataset size, we introduce a new track around open-vocabulary emotion recognition. The main consideration for this track is that existing datasets often fix the label space and use majority voting to enhance annotator consistency, but this process may limit the model’s ability to describe subtle emotions. In this track, we encourage participants to generate any number of labels in any category, aiming to describe the character’s emotional state as accurately as possible. Our baseline is based on MERTools [^3] https://github.com/zeroQiaoba/MERTools and the code is available at: https://github.com/zeroQiaoba/MERTools/tree/master/MER2024.
@@ 1. Introduction
Multimodal emotion recognition plays an important role in human-computer interaction. With the development of deep learning, researchers have made significant progress in this task. However, it still cannot be well addressed, and the emotion recognition performance still cannot meet the requirements of practical applications (lian2024merbench, ). To promote the development of this field, last year, we launched MER2023 (lian2023mer, ), focusing on three interesting topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue the latter two tracks and add a new track around open-vocabulary emotion recognition, aiming to describe the character’s emotional state as accurately as possible.
@@
Furthermore, emotions are subjective and ambiguous. To increase annotation consistency, existing datasets typically limit the label space to a few discrete labels, employ multiple annotators, and use majority voting to select the most likely label. However, this process may cause some correct but non-candidate or non-majority labels to be ignored, resulting in inaccurate annotations. To this end, we introduce a new track around open-vocabulary emotion recognition. We encourage participants to generate any number of labels in any category, trying to describe emotions accurately.
To sum up, MER2024 contains three tracks: MER-SEMI, MER-NOISE, and MER-OV. In MER-SEMI, we encourage participants to use unlabeled data for better performance; in MER-NOISE, we focus on the systems’ noise robustness; in MER-OV, we ask participants to describe the character’s emotional state as accurately as possible. This challenge provides participants with a platform to fairly compare the performance of different methods. In the remainder of this paper, we introduce the dataset, baseline systems, evaluation metrics, and experimental results for each track.
@@ 2.1. Data Collection
Video Cutting. To simulate real scenarios, we collect raw video from movies and TV series. However, the raw video is often long and have multiple characters. It is impossible to assign an emotion label to the entire video. Therefore, we need to split the raw video into video clips. In this process, we require that the video clips should contain relatively complete content from the same speaker, which is beneficial to understanding the emotional state of the character. Considering that the timestamps in subtitles are generally accurate, for videos with subtitles, we directly use these timestamps for video segmentation. For videos without subtitles, we use the voice activity detection toolkit, Silero VAD [^4] https://github.com/snakers4/silero-vad, to segment the video and the speaker identification toolkit, Deep Speaker [^5] https://github.com/philipperemy/deep-speaker, to merge the consecutive clips if they are likely to be from the same speaker.
Video Filtering. In the previous step, we aim to generate video clips containing relatively complete content from the same speaker. However, it only guarantees that the audio is from the same character, not that there is one character in the visual frames. Therefore, we further filter these video clips. Specifically, we use the face detection toolkit, YuNet [^6] https://github.com/ShiqiYu/libfacedetection, to ensure that most frames contain only one face. Then, we use the face recognition toolkit, face.evoLVe [^7] https://github.com/ZhaoJ9014/face.evoLVe, to ensure that these faces belong to the same person. Additionally, the length of video clips is also important. Video clips that are too short generally cannot express the complete meaning of the character; video clips that are too long may contain emotional changes, making it difficult to accurately describe the character’s emotional state with one label. Therefore, we perform an additional filtering operation, only selecting video clips between 2$\sim$16 seconds.
@@ 2.3. MER-OV
MER-OV focuses on open-vocabulary emotion recognition. Unlike MER-SEMI and MER-NOISE which only predict the most likely emotion within a fixed label space, MER-OV needs to predict any number of emotions in any category, trying to describe the character’s emotional state as accurately as possible. Previously, we made an initial attempt on this task (lian2023explainable, ). Specifically, we first annotate emotion-related acoustic and visual clues. Then, we rely on the LLMs’ reasoning ability and use these clues to disambiguate video subtitles. This process can generate descriptions containing more subtle emotional labels. Finally, we use ChatGPT to extract all labels using the following prompt:
@@
Table 2. Prompt for generating emotion-related descriptions using MLLMs.
@@ 2.4. Challenge Protocol
To download the dataset, participants should fill out an End User License Agreement (EULA) [^8] https://drive.google.com/file/d/1I0ZuPzL96W4Ow3KdF3N6D-mDhS5qJc-x. It requires participants to use this dataset only for academic research and not to edit or upload it to the Internet. For MER-SEMI and MER-NOISE, each team should submit the most likely discrete label among the 6 candidate labels (i.e., worried, happy, neutral, angry, surprised, and sad). For MER-OV, each team can submit any number of labels in any category. For all tracks, participants should predict results for all 115,595 unlabeled data, although we only evaluate a small subset. This requires participants to focus on the generalization ability and develop systems that do not require a large amount of inference time. Additionally, it is worth noting that participants cannot manually annotate the samples in MER2024. We will ask them to submit their code for further checking. Finally, each team should submit a paper describing their method, otherwise, the submitted results will not be included in the final leaderboard.
@@ 3.2. MER-OV
Open-vocabulary emotion recognition is a new task that lacks large-scale training data. Considering that MLLMs can handle various multimodal tasks, we choose MLLMs as the baselines. Meanwhile, temporal information is important for understanding the character’s emotions. Therefore, we select MLLMs that support at least audio or video input, including video LLMs (such as VideoChat (li2023videochat, )), audio LLMs (such as SALMONN (tang2023salmonn, )), and audio-video LLMs (such as PandaGPT (su2023pandagpt, )). In practice, we first use the prompt in Table 2 to generate emotion-relevant descriptions that contain multi-faceted and multi-modal clues. During inference, we add subtitles as additional input. Then, we use the prompt in Section 2.3 to extract all emotion labels in an open-vocabulary manner.
@@ 3.3. Implementation Details
Table 3. Unimodal results. In addition to five-fold cross-validation results on Train&\&&Val, we also report results on MER-SEMI and MER-NOISE. These features are ordered by the average WAF score on Train&\&&Val, MER-SEMI, and MER-NOISE. The values in gray columns are used as the final ranking metric for these tracks.
@@
Table 4. Multimodal results. We report both bimodal and trimodal results. Here, “A”, “V”, “T” represent acoustic, visual, and textual modalities, respectively. “TopN” means that we select the top-N features for each modality. Their ranking is based on the average WAF in Table 3.
@@ 4.1. MER-SEMI and MER-NOISE
Table 4 reports multimodal results. In this table, we choose “TopN” features and use the attention mechanism for multimodal fusion. “TopN” means that we select the top-$N$ features for each modality, and their ranking is based on the average WAF in Table 3. Experimental results show that different modality combinations prefer distinct $N$. Therefore, we should adjust this parameter carefully. Meanwhile, we can achieve the best result when considering trimodal information, reflecting the effectiveness of multimodal fusion. In terms of WAF scores, our baseline system can reach 86.73 in MER-SEMI and 79.62 in MER-NOISE.
Table 5. Performance of different methods on MER-OV (reported in our previous work (lian2023explainable, )). The values in the gray column are used for the final ranking. Here, “L”, “V”, and “A” indicate whether lexical, visual, and acoustic information is used in the emotion recognition process.
@@ 5. Conclusions
This paper introduces a multimodal emotion recognition challenge, MER2024, which consists of three tracks: MER-SEMI, MER-NOISE, and MER-OV. This challenge is an extended version of MER2023 with more labeled and unlabeled samples. Meanwhile, it introduces a new task called open-vocabulary emotion recognition, which aims to generate any number of labels in any category to describe the character’s emotional state as accurately as possible. In this paper, we introduce the datasets, establish the baselines, define the evaluation metrics, and report the baseline results for three tracks. Meanwhile, we open-source the code to ensure reproducibility. We hope that the MER series of challenges can provide researchers with a platform to evaluate their emotion recognition systems and prompt the application of this technology in real-life scenarios.

### v2

@@ Abstract.
Multimodal emotion recognition is an important research topic in artificial intelligence. Over the past few decades, researchers have made remarkable progress by increasing dataset size and building more effective architectures. However, due to various reasons (such as complex environments and inaccurate labels), current systems still cannot meet the demands of practical applications. Therefore, we plan to organize a series of challenges around emotion recognition to further promote the development of this field. Last year, we launched MER2023111http://merchallenge.cn/mer2023, focusing on three topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue to organize MER2024222https://zeroqiaoba.github.io/MER2024-website. In addition to expanding the dataset size, we introduce a new track around open-vocabulary emotion recognition. The main consideration for this track is that existing datasets often fix the label space and use majority voting to enhance annotator consistency, but this process may limit the model’s ability to describe subtle emotions. In this track, we encourage participants to generate any number of labels in any category, aiming to describe the emotional state as accurately as possible. Our baseline is based on MERTools [^3] https://github.com/zeroQiaoba/MERTools and the code is available at: https://github.com/zeroQiaoba/MERTools/tree/master/MER2024.
@@ 1. Introduction
Multimodal emotion recognition plays an important role in human-computer interaction. With the development of deep learning, researchers have made significant progress in this task. However, it still cannot be well addressed, and the emotion recognition performance still cannot meet the requirements of practical applications (lian2024merbench, ). To promote the development of this field, last year, we launched MER2023 (lian2023mer, ), focusing on three interesting topics: multi-label learning, noise robustness, and semi-supervised learning. This year, we continue the latter two tracks and add a new track around open-vocabulary emotion recognition, aiming to describe the emotional state as accurately as possible.
@@
Furthermore, emotions are subjective and ambiguous. To increase annotation consistency, existing datasets typically limit the label space to a few discrete labels. Then, they employ multiple annotators and use majority voting to select the most likely label. However, this process may cause some correct but non-candidate or non-majority labels to be ignored, resulting in inaccurate annotations. To this end, we introduce a new track around open-vocabulary emotion recognition. We encourage participants to generate any number of labels in any category, trying to describe emotions accurately.
To sum up, MER2024 contains three tracks: MER-SEMI, MER-NOISE, and MER-OV. In MER-SEMI, we encourage participants to use unlabeled data for better performance; in MER-NOISE, we focus on the systems’ noise robustness; in MER-OV, we ask participants to describe the emotional state as accurately as possible. This challenge provides participants with a platform to fairly compare the performance of different methods. In the remainder of this paper, we introduce the dataset, baseline systems, evaluation metrics, and experimental results for each track.
@@ 2.1. Data Collection
Video Cutting. To simulate real scenarios, we collect raw video from movies and TV series. However, the raw video is often long and have multiple characters. It is impossible to assign an emotion label to the entire video. Therefore, we need to split the raw video into video clips. In this process, we require that the video clips should contain relatively complete content from the same character. Considering that the timestamps in subtitles are generally accurate, for videos with subtitles, we directly use these timestamps for video segmentation. For videos without subtitles, we use the voice activity detection toolkit, Silero VAD [^4] https://github.com/snakers4/silero-vad, to segment the video and the speaker identification toolkit, Deep Speaker [^5] https://github.com/philipperemy/deep-speaker, to merge the consecutive clips if they are likely to be from the same speaker.
Video Filtering. In the previous step, we aim to generate video clips containing relatively complete content from the same speaker. However, it only guarantees that the audio is from the same character, not that there is one character in the visual frames. Therefore, we further filter these video clips. Specifically, we use the face detection toolkit, YuNet [^6] https://github.com/ShiqiYu/libfacedetection, to ensure that most frames contain only one face. Then, we use the face recognition toolkit, face.evoLVe [^7] https://github.com/ZhaoJ9014/face.evoLVe, to ensure that these faces belong to the same person. Additionally, the length of video clips is also important. Video clips that are too short generally cannot express the complete meaning; video clips that are too long may contain emotional changes, making it difficult to accurately describe the emotional state with one label. Therefore, we perform an additional filtering operation, only selecting video clips between 2$\sim$16 seconds.
@@ 2.3. MER-OV
MER-OV focuses on open-vocabulary emotion recognition. Unlike MER-SEMI and MER-NOISE which only predict the most likely emotion within a fixed label space, MER-OV needs to predict any number of emotions in any category, trying to describe the emotional state as accurately as possible. Previously, we made an initial attempt on this task (lian2023explainable, ). Specifically, we first annotate emotion-related acoustic and visual clues. Then, we rely on the LLMs’ reasoning ability and use these clues to disambiguate video subtitles. This process can generate descriptions containing more subtle emotional labels. Finally, we use ChatGPT to extract all labels using the following prompt:
@@
Table 2. Prompt for generating emotion-related descriptions using MLLMs (drawn from our previous work (lian2023explainable, )).
@@ 2.4. Challenge Protocol
To download the dataset, participants should fill out an End User License Agreement (EULA) [^8] https://drive.google.com/file/d/1cXNfKHyJzVXg_7kWSf_nVKtsxIZVa517/view?usp=sharing. It requires participants to use this dataset only for academic research and not to edit or upload it to the Internet. For MER-SEMI and MER-NOISE, each team should submit the most likely discrete label among the 6 candidate labels (i.e., worried, happy, neutral, angry, surprised, and sad). For MER-OV, each team can submit any number of labels in any category. For all tracks, participants should predict results for all 115,595 unlabeled data, although we only evaluate a small subset. This requires participants to focus on the generalization ability and develop systems that do not require a large amount of inference time. Additionally, it is worth noting that participants cannot manually annotate the samples in MER2024. We will ask them to submit their code for further checking. Finally, each team should submit a paper describing their method, otherwise, the submitted results will not be included in the final leaderboard.
@@ 3.2. MER-OV
Open-vocabulary emotion recognition is a new task that lacks large-scale training data. Considering that MLLMs can handle various multimodal tasks, we choose MLLMs as the baselines. Meanwhile, temporal information is important for understanding the emotions. Therefore, we select MLLMs that support at least audio or video input, including video LLMs (such as VideoChat (li2023videochat, )), audio LLMs (such as SALMONN (tang2023salmonn, )), and audio-video LLMs (such as PandaGPT (su2023pandagpt, )). In practice, we first use the prompt in Table 2 to generate emotion-relevant descriptions that contain multi-faceted and multi-modal clues. During inference, we add subtitles as additional input. Then, we use the prompt in Section 2.3 to extract all emotion labels in an open-vocabulary manner.
@@ 3.3. Implementation Details
Table 3. Unimodal results (%). In addition to five-fold cross-validation results on Train&\&&Val, we also report results on MER-SEMI and MER-NOISE. These features are ordered by the average WAF score on Train&\&&Val, MER-SEMI, and MER-NOISE. The values in gray columns are used as the final ranking metric for these tracks.
@@
Table 4. Multimodal results (%). We report both bimodal and trimodal results. Here, “A”, “V”, “T” represent acoustic, visual, and textual modalities, respectively. “TopN” means that we select the top-N features for each modality. Their ranking is based on the average WAF in Table 3.
@@ 4.1. MER-SEMI and MER-NOISE
Table 4 reports multimodal results. In this table, we choose “TopN” features and use the attention mechanism for multimodal fusion. “TopN” means that we select the top-$N$ features for each modality, and their ranking is based on the average WAF in Table 3. Experimental results show that different modality combinations prefer distinct $N$. Therefore, we should adjust this parameter carefully. Meanwhile, we can achieve the best result when considering trimodal information, reflecting the effectiveness of multimodal fusion. In terms of WAF scores, our baseline system can reach 86.73% in MER-SEMI and 79.62% in MER-NOISE.
Table 5. Performance of different methods on MER-OV (%). These results are drawn from our previous work (lian2023explainable, ). The values in the gray column are used for the final ranking. Here, “L”, “V”, and “A” indicate whether lexical, visual, and acoustic information is used in the emotion recognition process.
@@ 5. Conclusions
This paper introduces a multimodal emotion recognition challenge, MER2024, which consists of three tracks: MER-SEMI, MER-NOISE, and MER-OV. This challenge is an extended version of MER2023 with more labeled and unlabeled samples. Meanwhile, it introduces a new task called open-vocabulary emotion recognition, which aims to generate any number of labels in any category to describe the emotional state as accurately as possible. In this paper, we introduce the datasets, establish the baselines, define the evaluation metrics, and report the baseline results for three tracks. Meanwhile, we open-source the code to ensure reproducibility. We hope that the MER series of challenges can provide researchers with a platform to evaluate their emotion recognition systems and prompt the application of this technology in real-life scenarios.